# Denoising with a Joint-Embedding Predictive Architecture

### v1

@@ Abstract
Joint-embedding predictive architectures (JEPAs) have shown substantial promise in self-supervised representation learning, yet their application in generative modeling remains underexplored. Conversely, diffusion models have demonstrated significant efficacy in modeling arbitrary probability distributions. In this paper, we introduce Denoising with a Joint-Embedding Predictive Architecture (D-JEPA), pioneering the integration of JEPA within generative modeling. By recognizing JEPA as a form of masked image modeling, we reinterpret it as a generalized next-token prediction strategy, facilitating data generation in an auto-regressive manner. Furthermore, we incorporate diffusion loss to model the per-token probability distribution, enabling data generation in a continuous space. We also adapt flow matching loss as an alternative to diffusion loss, thereby enhancing the flexibility of D-JEPA. Empirically, with increased GFLOPs, D-JEPA consistently achieves lower FID scores with fewer training epochs, indicating its good scalability. Our base, large, and huge models outperform all previous generative models across all scales on class-conditional ImageNet benchmarks. Beyond image generation, D-JEPA is well-suited for other continuous data modeling, including video and audio.
@@ 1 Introduction
Recently, diffusion models (Ho et al., 2020; Song et al., 2020; Nichol & Dhariwal, 2021) have emerged as a dominant force in generative modeling, surpassing generative adversarial networks (GANs) (Goodfellow et al., 2014; Wang et al., 2017; Goodfellow et al., 2020) with their superior performance in producing high-fidelity images. Concurrently, significant advancements have been made in self-supervised representation learning, particularly through masked image modeling (Xie et al., 2022; Baevski et al., 2022; Assran et al., 2023; He et al., 2022a) and invariance-based techniques (Assran et al., 2022; Caron et al., 2021; Chen & He, 2021; Grill et al., 2020; Zbontar et al., 2021). However, generative modeling and representation learning development has largely progressed in isolation, with limited cross-fertilization (Li et al., 2023).
@@
Recent efforts have sought to integrate diffusion models within an autoregressive framework, such as MAR (Li et al., 2024) and Transfusion (Zhou et al., 2024). These models, however, primarily focus on enhancing generative model performance from the perspective of generation alone, often overlooking advancements in representation learning. Conversely, within the realm of representation learning, researchers tend to emphasize model performance on downstream tasks  (Assran et al., 2023), neglecting the potential for these models to excel in generative modeling. To bridge this gap, we propose a novel framework, termed “Denoising with a Joint-Embedding Predictive Architecture” (D-JEPA), that effectively leverages next-token prediction from a representation learning perspective. Specifically, we adopt the joint-embedding predictive architecture as the core of our design, enabling us to reframe masked image modeling methods into a generalized next-token prediction strategy, thus facilitating autoregressive data generation. Additionally, we incorporate a diffusion loss (or flow matching loss) to model the per-token probability distribution, allowing for data generation in a continuous space.
@@
Specifically, D-JEPA consists of three identical visual transformer backbones (Dosovitskiy et al., 2020): a context encoder $\phi$, a target encoder $\bar{\phi}$, and a feature predictor $\gamma$. It employs two loss functions: diffusion loss ($\mathcal{L}_{d}$) and prediction loss ($\mathcal{L}_{p}$), both involving a multi-layer perceptron (MLP) network. The denoising MLP, used in diffusion loss, is applied to each masked token $x_{i}$, modeling the conditional probability distribution $p(x_{i}\mid z_{i})$ and evaluating the predicted latent variable $z_{i}$ generated by the feature predictor $\gamma$. The prediction loss ($\mathcal{L}_{p}$), using a smooth loss function $l_{1}$, facilitates the latent variables $z_{i}$ to capture high-level semantic information.
@@
Empirically, D-JEPA exhibits a consistent decrease in FID scores as GFLOPs increase, achieving these results with fewer training epochs, indicating its good scalability. Our base, large, and huge variants outperform all previously generative models across different scales on class-conditional ImageNet benchmarks. In our comprehensive experiments, we further advanced the field of generative models by substituting the diffusion loss with flow matching loss (Lipman et al., 2022). This substitution enables D-JEPA to excel in image generation and modeling other types of continuous data, including video and audio. Moreover, D-JEPA exhibits remarkable flexibility, extending seamlessly to text-conditioned and multi-modal generations.
@@
Our contributions are threefold. Firstly, we successfully integrate advancements in representation learning into generative modeling, achieving state-of-the-art performance on the ImageNet benchmark. Secondly, we validate the effectiveness of the generalized next-token prediction approach in generative modeling, demonstrating superior scalability, which establishes a robust foundation for developing more powerful text-conditioned models. Lastly, D-JEPA exhibits impressive extensibility, directly applicable to generate images, videos, audio, and more, providing theoretical support for constructing unified multi-modal models.
@@ 2 Background
However, the issue of representation collapse is effectively mitigated in D-JEPA due to the introduction of an additional diffusion loss for the predicted embeddings. The diffusion loss does not reside in the same energy landscape as JEPAs, thus effectively avoiding model collapse. Consequently, similar to the approach in LeCun (2022), we can employ two identical networks as the context and target encoders. The parameters of the target encoder are efficiently updated through an exponential moving average method, significantly enhancing training efficiency and reducing the number of parameters that need to be learned. Refer to Sec. 3.2 for more details.
@@
Our D-JEPA model extends these innovations by integrating the masked image modeling approach as a generalized next-token prediction strategy to facilitate an autoregressive generation manner. We can flexibly adjust the number of tokens predicted in each step by controlling the auto-regressive steps. Contrary to the intuition formed by existing methods (Li et al., 2024; 2023), the best performance under the D-JEPA architecture is not achieved when the auto-regressive steps are maximized, i.e., when the generalized next-token prediction predicts only one token at a time. Moreover, as the scale of the model parameters increases, the number of auto-regressive steps required to achieve the best sampling results decreases accordingly. This fully demonstrates that converting the masked image modeling method into a next-token prediction strategy is not only feasible but also highly effective. Refer to the Experiment part in Sec. 4 for more analysis.
@@ 3.1 Tokenization and Masking
Although images are directly segmented into patches on raw pixels, which is also the standard practice in visual transformers (Dosovitskiy et al., 2020), recent generative models (Li et al., 2024; 2023; Rombach et al., 2022) tend to first encode images into the latent space using a VAE before performing patchification, arguing that this significantly reduces the training cost of generative models and accelerates the sampling speed. However, we found that this approach noticeably degrades D-JEPA’s performance in representation tasks. [^1] More analysis about D-JEPA’s performance in representation tasks have been listed in Appendix E.2. Since the primary goal of this work is to ensure that D-JEPA functions as an outstanding generative model, we adopted the latent space modeling strategy initially proposed by Rombach et al. (2022). (It is important to note that training in latent space is not a necessity for D-JEPA; it can be trained in raw space and still achieve excellent results.)
@@ 3.2 Representation Learning with JEPAs
Earlier masked image modeling methods, such as MAE (He et al., 2022a), directly use a decoder network to predict the missing raw pixels. This approach gives these models some outpainting capabilities (i.e., completing missing parts of an image), but they still cannot directly generate high-quality images. Subsequent works, such as DiffMAE (Wei et al., 2023), attempted to leverage diffusion models within MAE to enhance the image generation capabilities of these models, but the results were limited. Moreover, this direct raw pixel prediction approach severely restricts the model’s representation ability, especially compared to models like CLIP (Radford et al., 2021), which use contrastive learning on massive datasets.
@@ 4.1 Image Synthetis
We present the Fréchet Inception Distance (FID) along with other metrics in Table 1 without classifier-free guidance. Two critical conclusions can be drawn from the results: State-of-the-Art Performance: D-JEPA achieves state-of-the-art FID and Inception Score (IS) across all model scales. In particular, D-JEPA-L, with only 687M parameters, surpasses all previous works with $\text{FID}=2.32,\text{IS}=233.5$. The larger model, D-JEPA-H, further improves the benchmark set by D-JEPA-L to $\text{FID}=2.04,\text{IS}=239.3$. Efficiency in Training: As the model scale increases, the required number of training epochs for D-JEPA decreases significantly. Both D-JEPA-L and D-JEPA-H achieve state-of-the-art performance with substantially fewer training epochs compared to other models. For instance, D-JEPA-L, with 687M parameters and 480 training epochs, outperforms the previous state-of-the-art model, MAR-H, which has 943M parameters and requires 800 training epochs.
It is well known that selecting an appropriate cfg scale can significantly enhance the quality of generated images. For D-JEPA, in addition to adjusting the cfg scale, it is also necessary to choose an appropriate temperature $\tau$ to ensure optimal generation quality. We employ a coarse-to-fine grid search to determine the optimal sampling parameters, as detailed in the Appendix C.2. The optimal sampling configurations are listed on Tab. 5.
@@
Table 2: System-level comparison on ImageNet $256\times 256$ conditional generation. To better showcase model performance, each model is evaluated with two different cfg scales (obtained by grid searching). One cfg scale provides the best FID score, while the other yields a better IS result at the cost of a negligible degradation in the FID score.
@@ 4.2 Comprehensive Experiments
In the Appendix, we further demonstrate the excellent inpainting and outpainting capabilities of the D-JEPA model (Appendix D.3). Beyond diffusion loss, we also designed a flow matching loss to construct D-JEPA (Appendix F). Experiments show that flow matching loss performs better on many tasks. Consequently, we extended our experiments based on flow matching loss to include text-to-image/audio generation and class-conditional video generation. These extended experiments fully demonstrate that D-JEPA is capable of generating various types of continuous data, not limited to images. Furthermore, we attempted to construct a multimodal generation model based on D-JEPA. The results indicate that D-JEPA can generate continuous data and be used to build multimodal understanding and generation models (Appendix G).
@@ Appendix A Related Work
Recent advances such as MAGE (Li et al., 2023) achieve high-quality representations and image generation. However, these methods are limited to image data, failing to empirically demonstrate mutual reinforcement between tasks. Our work uniquely verifies the mutual benefits of representation learning and generative modeling through empirical evidence, demonstrating how these processes can be synergistically combined for superior performance across different types of continuous data.
@@ D.1 Full comparison on ImageNet
D.2 Training curve

### v2

@@ Abstract
Joint-embedding predictive architectures (JEPAs) have shown substantial promise in self-supervised representation learning, yet their application in generative modeling remains underexplored. Conversely, diffusion models have demonstrated significant efficacy in modeling arbitrary probability distributions. In this paper, we introduce Denoising with a Joint-Embedding Predictive Architecture (D-JEPA), pioneering the integration of JEPA within generative modeling. By recognizing JEPA as a form of masked image modeling, we reinterpret it as a generalized next-token prediction strategy, facilitating data generation in an auto-regressive manner. Furthermore, we incorporate diffusion loss to model the per-token probability distribution, enabling data generation in a continuous space. We also adapt flow matching loss as an alternative to diffusion loss, thereby enhancing the flexibility of D-JEPA. Empirically, with increased GFLOPs, D-JEPA consistently achieves lower FID scores with fewer training epochs, indicating its good scalability. Our base, large, and huge models outperform all previous generative models across all scales on ImageNet conditional generation benchmarks. Beyond image generation, D-JEPA is well-suited for other continuous data modeling, including video and audio.
@@ 1 Introduction
Recently, diffusion models (Ho et al., 2020; Song et al., 2020; Nichol & Dhariwal, 2021) have emerged as a dominant force in generative modeling, surpassing generative adversarial networks (GANs) (Goodfellow et al., 2014; Wang et al., 2017; Goodfellow et al., 2020) with their superior performance in producing high-fidelity images. Concurrently, significant advancements have been made in self-supervised representation learning, particularly through masked image modeling (Xie et al., 2022; Baevski et al., 2022; Assran et al., 2023; He et al., 2022a) and invariance-based techniques (Assran et al., 2022; Caron et al., 2021; Chen & He, 2021; Grill et al., 2020; Zbontar et al., 2021). However, generative modeling has yet to directly benefit from the advancements in representation learning (Li et al., 2023).
@@
Recent efforts have sought to integrate diffusion models within an autoregressive framework, such as MAR (Li et al., 2024) and Transfusion (Zhou et al., 2024). These models, however, primarily focus on enhancing generative model performance from the perspective of generation alone, often overlooking the potential advancements that can be brought by representation learning. Conversely, within the realm of representation learning, researchers tend to emphasize model performance on downstream tasks  (Assran et al., 2023), neglecting the potential for these models to excel in generative modeling. To overcome this challenge, we propose a novel framework, termed “Denoising with a Joint-Embedding Predictive Architecture” (D-JEPA), that effectively leverages next-token prediction from a representation learning perspective to generative modeling. Specifically, we adopt the joint-embedding predictive architecture as the core of our design, enabling us to reframe masked image modeling methods into a generalized next-token prediction strategy, thus facilitating autoregressive data generation. Additionally, we incorporate a diffusion loss (or flow matching loss) to model the per-token probability distribution, allowing for data generation in a continuous space.
@@
D-JEPA consists of three identical visual transformer backbones (Dosovitskiy et al., 2020): a context encoder $\phi$, a target encoder $\bar{\phi}$, and a feature predictor $\gamma$. It employs two loss functions: diffusion loss ($\mathcal{L}_{d}$) and prediction loss ($\mathcal{L}_{p}$), both involving a multi-layer perceptron (MLP) network. The denoising MLP, used in diffusion loss, is applied to each masked token $x_{i}$, modeling the conditional probability distribution $p(x_{i}\mid z_{i})$ and evaluating the predicted latent variable $z_{i}$ generated by the feature predictor $\gamma$. The prediction loss ($\mathcal{L}_{p}$), using a smooth loss function $l_{1}$, facilitates the latent variables $z_{i}$ to capture high-level semantic information.
@@
Empirically, D-JEPA exhibits a consistent decrease in FID scores as GFLOPs increase, achieving these results with fewer training epochs, indicating its good scalability. Our base, large, and huge variants outperform all previously generative models across different scales on ImageNet conditional generation benchmarks. In our comprehensive experiments, we further advanced the field of generative models by substituting the diffusion loss with flow matching loss (Lipman et al., 2022). This substitution enables D-JEPA to excel in image generation and modeling other types of continuous data, including video and audio. Moreover, D-JEPA exhibits remarkable flexibility, extending seamlessly to text-conditioned and multi-modal generations.
@@
Our contributions are threefold. Firstly, we successfully integrate advancements in representation learning into generative modeling, achieving state-of-the-art performance on the ImageNet conditional generation benchmarks. Secondly, we validate the effectiveness of the generalized next-token prediction approach in generative modeling, demonstrating superior scalability, which establishes a robust foundation for developing more powerful text-conditioned models. Lastly, D-JEPA exhibits impressive extensibility, directly applicable to generate images, videos, audio, and more, providing theoretical support for constructing unified multi-modal generation models for continuous data.
@@ 2 Background
However, the issue of representation collapse is effectively mitigated in D-JEPA due to the introduction of an additional diffusion loss for the predicted embeddings. The diffusion loss does not reside in the same energy landscape as JEPAs, thus effectively avoiding model collapse. Consequently, similar to the approach in LeCun (2022), we can employ two identical networks as the context and target encoders. The parameters of the target encoder are efficiently updated through an exponential moving average method, significantly enhancing training efficiency and reducing the number of parameters that need to be learned. Please refer to Sec. 3.2 for more details.
@@
Our D-JEPA model extends these innovations by integrating the masked image modeling approach as a generalized next-token prediction strategy to facilitate an autoregressive generation manner. We can flexibly adjust the number of tokens predicted in each step by controlling the auto-regressive steps. Contrary to the intuition formed by existing methods (Li et al., 2024; 2023), the best performance under the D-JEPA architecture is not achieved when the auto-regressive steps are maximized, i.e., when the generalized next-token prediction predicts only one token at a time. Moreover, as the scale of the model parameters increases, the number of auto-regressive steps required to achieve the best sampling results decreases accordingly. This fully demonstrates that converting the masked image modeling method into a next-token prediction strategy is not only feasible but also highly effective. Please refer to the experiment part in Sec. 4 for more analysis.
@@ 3.1 Tokenization and Masking
Although images are directly segmented into patches on raw pixels, which is also the standard practice in visual transformers (Dosovitskiy et al., 2020), recent generative models (Li et al., 2024; 2023; Rombach et al., 2022) tend to first encode images into the latent space using a VAE before performing patchification, arguing that this significantly reduces the training cost of generative models and accelerates the sampling speed. However, we found that this approach noticeably degrades D-JEPA’s performance in representation tasks. Since the primary goal of this work is to ensure that D-JEPA functions as an outstanding generative model, we adopted the latent space modeling strategy initially proposed by Rombach et al. (2022). (It is important to note that training in latent space is not a necessity for D-JEPA; it can be trained in raw space and still achieve excellent results. The experiments conducted in raw pixel space are detailed in Appendix E.2.)
@@ 3.2 Representation Learning with JEPAs
Earlier masked image modeling methods, such as MAE (He et al., 2022a), directly use a decoder network to predict the missing raw pixels. This approach gives these models some outpainting capabilities (i.e., completing missing parts of an image), but they still cannot directly generate high-quality images. Subsequent works, such as DiffMAE (Wei et al., 2023), attempted to leverage diffusion models within MAE (He et al., 2022a) to enhance the image generation capabilities of these models, but the results were limited. Moreover, this direct raw pixel prediction approach severely restricts the model’s representation ability, especially compared to models like CLIP (Radford et al., 2021), which use contrastive learning on massive datasets.
@@ 4.1 Image Synthetis
We present the Fréchet Inception Distance (FID) along with other metrics in Tab. 1 without classifier-free guidance. Two critical conclusions can be drawn from the results: State-of-the-Art Performance: D-JEPA achieves state-of-the-art FID and Inception Score (IS) across all model scales. In particular, D-JEPA-L, with only 687M parameters, surpasses all previous works across all scales with $\text{FID}=2.32,\text{IS}=233.5$. The larger model, D-JEPA-H, further improves the benchmark set by D-JEPA-L to $\text{FID}=2.04,\text{IS}=239.3$. Efficiency in Training: As the model scale increases, the required number of training epochs for D-JEPA decreases significantly. Both D-JEPA-L and D-JEPA-H achieve state-of-the-art performance with substantially fewer training epochs compared to other models. For instance, D-JEPA-L, with 687M parameters and 480 training epochs, outperforms the previous state-of-the-art model, MAR-H, which has 943M parameters and requires 800 training epochs. It is worth noting that D-JEPA-B achieved similar performance to MAR-B (FID=3.50 vs. 3.48) with only 600 training epochs compared to MAR-B’s 800 epochs. For a more detailed analysis, please refer to the ablation study on Appendix C.3.
It is well known that selecting an appropriate cfg scale (Ho & Salimans, 2021) can significantly enhance the quality of generated images. For D-JEPA, in addition to adjusting the cfg scale, it is also necessary to choose an appropriate temperature $\tau$ to ensure optimal generation quality. We employ a coarse-to-fine grid search to determine the optimal sampling parameters, as detailed in the Appendix C.2. The optimal sampling configurations are listed on Tab. 5.
@@
Table 2: System-level comparison on ImageNet $256\times 256$ conditional generation.
@@ 4.2 Comprehensive Experiments
In Appendix  D.4, we further demonstrate the excellent inpainting and outpainting capabilities of the D-JEPA model. Beyond diffusion loss, we also designed a flow matching loss to construct a more powerful D-JEPA (Appendix F). Experiments show that flow matching loss performs better on many tasks. Consequently, we extended our experiments based on flow matching loss to include text-to-image/audio generation and class-conditional video generation. These extended experiments fully demonstrate that D-JEPA is capable of generating various types of continuous data, not limited to images. Furthermore, we attempted to construct a multimodal generation model based on D-JEPA. The results indicate that D-JEPA can generate continuous data and be used to build multimodal understanding and generation models (Appendix G).
@@ Appendix A Related Work
Recent advances such as MAGE (Li et al., 2023) achieve high-quality representations and image generation. However, these methods are limited to image data, and have not been validated on other types of continuous data such as speech and video. Our work uniquely demonstrates the significant impact of representation learning across image, audio, and video generation, highlighting its versatility and effectiveness in these domains.
@@ C.2 Abalation on auto-regressive steps
C.3 Ablation on JEPA loss
We conducted ablation experiments on the D-JEPA-B model. The results of MAR-B serve as the baseline (i.e., the results without the JEPA Loss), because MAR-B and D-JEPA-B have identical model structures and experimental parameters, making them a comparable baseline.
From Fig. 6, we can see that D-JEPA-B achieves the performance of MAR-B at 600 epochs. Additionally, from Tab. 1 and Tab. 2, we observe that D-JEPA-B significantly outperforms MAR-B at 1400 epochs. For a more intuitive comparison, we provide some key data points without using classifier-free guidance in Tab. 6, which we believe better reflects the true scenario.
Model #params Epoch FID(w/o cfg) IS(w/o cfg) FID(w/ cfg) IS(w/ cfg)
MAR-B 208M 800 3.48 192.4 2.31 281.7
D-JEPA-B 208M+4M 600 3.50 189.2 2.25 279.8
800 3.43 195.3 2.01 280.1
1400 3.40 197.1 1.87 282.5
Table 6: Ablation on JEPA loss. The additional 4M parameters are introduced by the MLP layers in JEPA loss. CFG is set as 3.1 for D-JEPA-B.
@@ D.1 Full comparison on ImageNet
D.2 Additional comparison on ImageNet 512$\times$512
w/o CFG w/ CFG
[…]
pixel-based
ADM (2021) 554M 23.24 58.1 7.72 172.7
VDM++ (2024) 2B 2.99 232.2 2.65 278.1
vector-quantized tokens
MaskGIT (2022) 227M 7.32 156.0 - -
MAGVIT-v2 (2023b) 307M 3.07 213.1 1.91 324.3
continuous-valued tokens
U-ViT-H/2-G (2022) 501M - - 4.05 263.8
DiT-XL/2 (2023) 675M 12.03 105.3 3.04 240.8
DiffiT (2025) - - - 2.67 252.1
GIVT (2025) 304M 8.35 - - -
EDM2-XXL (2024) 1.5B 1.91 - 1.81 -
MAR-L (2024) 481M 2.74 205.2 1.73 279.9
D-JEPA-L 687M 1.89 239.1 1.62 335.2
Table 8: System-level comparison on ImageNet 512$\times$512 conditional generation. D-JEPA-L is trained for 480 epochs.
Following previous works, we also report results on ImageNet at a resolution of 512$\times$512, compared with leading systems (see Tab. 8). For simplicity, we use the KL-16 tokenizer, which results in a sequence length of 32$\times$32 for a 512$\times$512 image. Other settings follow the D-JEPA-L configuration. Our method achieves a Fréchet Inception Distance (FID) of 1.89 without Classifier-Free Guidance (CFG) and 1.62 with CFG. Our results surpass those of previous systems. Due to limited resources, we have not trained the larger D-JEPA-H model on ImageNet 512$\times$512, which is expected to yield even better results.
D.3 Training curve
@@ E.2 Image classification
Continuous tokens Raw pixels
[…]
D-JEPA-B w/o 3.40 197.1 3.45 195.2
D-JEPA-B 3.1 1.87 282.5 1.89 281.3
D-JEPA-B 4.1 2.08 320.9 2.11 320.1
D-JEPA-L w/o 2.32 233.5 2.30 235.0
D-JEPA-L 3.0 1.58 303.1 1.61 300.7
D-JEPA-L 3.9 1.65 327.2 1.66 330.3
Table 11: Comparison of generative performance of D-JEPA trained on continuous tokens and raw pixels.
Generative performance on raw pixel space and semantic token space.
We evaluated the generative performance of D-JEPA on both the raw pixel space and the semantic token space. As shown in Tab. 11, D-JEPA achieves excellent results in both spaces. Training in the raw pixel space requires significantly more computational resources, which is why we conducted experiments with D-JEPA-B and D-JEPA-L, omitting D-JEPA-H. However, based on the results, we believe that D-JEPA-H can also perform well in the raw space, albeit at the cost of greater computational resources.